# Phase transition on the fluctuation of the structure of random walk ranges

### v1

@@ 6.1. Lower bounds for the Cross Term
We will first start by obtaining a lower bound of $\mathbb{E}[E_{n}].$ Our heuristic argument is the following: if there is an intersection between $S[3n/2,2n]$ and $S[0,n]$, then this will form a shortcut in the graph distance between the points $S_{2n}$ and $S_{0}$. The amount that the graph distance would reduce via this shortcut will be related to the graph distance between $S_{n}$ and $S_{2n}$, which by a law of large numbers should be $O\left(\frac{n}{(\log n)^{1/2}}\right)$. Furthermore, the probability of the intersection of $S[3n/2,2n]$ and $S[0,n]$ is approximately $O((\log n)^{-1})$. Thus, one would heuristically argue that $\mathbb{E}[E_{n}]=\frac{n}{(\log n)^{3/2\pm o(1)}}$.
However, the exact value of the reduction of the graph distance is a complicated function of the loop intersection structure of the two walks. Thus, knowing that there is a long-range intersection may not be enough the deduce the reduction of the graph distance. Furthermore, to get the exact order $\frac{n}{(\log n)^{3/2}}$, that we desire, then we would need to argue that there is exact independence between the event that there is an intersection between $S[3n/2,2n]$ and $S[0,n]$ and the event that the graph distance between $S_{n}$ and $S_{2n}$ is $O\left(\frac{n}{(\log n)^{1/2}}\right)$. The following lemma will address these issues by finding appropriate events that will allow us to constrain the geometry of the random walk range.
@@
Let $E_{n}$ denote the cross term that appears when considering the number of cut points, the effective resistance or the graph distance. We have $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$.
@@
In [21, Section 4], he has already shown $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$ for the effective resistance. We obtain the same result by a slightly simpler proof which derives intersection estimates by using the capacity of the random walk.
@@
$A_{5}=\{\tilde{S}_{i}\leq C_{1}\sqrt{n},\quad\forall i\in[n-n/(\log n)^{\epsilon},n]\}$
noting $\mathbb{E}[{\rm Cap}(\tilde{S}[0,n])]\sim\frac{\pi^{2}}{8}\frac{n}{\log n}$ (e.g., see [6]). In the above equation, we use the notation $\tilde{X}[a,b]$ to denote the graph distance between the points $\tilde{S}_{a}$ and $\tilde{S}_{b}$ along the graph $\tilde{S}[a,b]$. By [11, Lemma 2.3], $\mathbb{P}(A_{1})=1-o(1)$. Similarly, the law of large numbers ensures that $\mathbb{P}(A_{2})$, $\mathbb{P}(A_{3})=1-o(1)$ noting [11, Theorem 1.2] and $\mathbb{P}(A_{4})=1-o(1)$ (see [6, Theorem 1.1]). By the local central limit theorem, we have that $\mathbb{P}(A_{5})=1-\epsilon(C_{1})$, where $\epsilon(C_{1})\to 0$ as $C_{1}\to\infty$. Let $\mathcal{Q}:=A_{1}\cap A_{2}\cap A_{3}\cap A_{4}\cap A_{5}$. From our previous estimates, we can take a union bound to deduce that $\mathbb{P}(\mathcal{Q})\geq 1-2\epsilon(C_{1})$, for $n$ sufficiently large.
Now, we need to define an event on the intersection of $\hat{S}$ with $\tilde{S}$. Let $\tau$ be defined to be the first time of intersection between $\hat{S}$ and $\tilde{S}[n-n/(\log n)^{\epsilon},n]$ after time $n-n/(\log n)^{\epsilon}$. That is, $\tau:=\inf\{i>n-n/(\log n)^{\epsilon}:\hat{S}_{i}\in\tilde{S}[n-n/(\log n)^{\epsilon},n]\}$. We define,
@@
On the union $A_{5}\cap B_{2}$, we can apply the local central limit theorem to deduce that $\left(\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\right)\geq cn^{-1}(\log n)^{-\epsilon}$, for some constant $c>0$ on $\mathcal{Q}$. Furthermore, the event $A_{5}$ ensures that ${\rm Cap}(\mathcal{S})\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}$ on $\mathcal{Q}$. Multiplying these together, we see that there is some $c>0$ such that $\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\tau<n/(\log n)^{\epsilon})>c/(\log n)^{1+2\epsilon}$ on $\mathcal{Q}$. This means that $\mathbb{P}(B_{3}|B_{2}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$. Returning to equation (6.1), we see that we see that we must have $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$.
Now, on the event $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$, we claim we have $E_{n}>n/(\log n)^{1/2+\epsilon/2}$. Due to the presence of the cut points from the event $A_{1}$ and the lower bound on the graph distance $\tilde{X}_{n/2}$ from $A_{2}$, we have that $\tilde{X}_{n}\geq\frac{1}{1+\epsilon}\mathbb{E}[X_{n/2}]-\frac{n}{(\log n)^{7/12}}$. By events $A_{3}$, $A_{1}$, and $B_{1}$, the graph distance between the points $\hat{S}_{n}$ and $\tilde{S}_{n}$, is at most $2n/(\log n)^{1/2+\epsilon/2}+2\frac{n}{(\log n)^{7/12}}$. Computing the difference, we now see that $E_{n}>n/(\log n)^{1/2+\epsilon/2}$ on $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$ as desired and this event happens with probability $c/(\log n)^{1+2\epsilon}$. By multiplying these two terms together, we see that $\mathbb{E}[E_{n}]>n/(\log n)^{3/2+\epsilon/2+2\epsilon}$. This proof also works even for the effective resistance or the number of cut points. ∎
@@
We claim both that $\Omega$ holds with probability $O\left(\frac{\log\log n}{\log n}\right)$ and that on the event $\Omega^{c}$, we must have that the cross term $E_{n}$ for number of cut points, the effective resistance, and the graph distance must by less than $\frac{n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$ by Lemma 6.1.
@@
Now, we will check that on $\Omega^{c}$, our cross term can be no more than $\frac{n}{(\log n)^{2}}$. Notice that on $\Omega^{c}$, there must be some cut point in the interval $\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]$ for both $\tilde{S}$ and $\hat{S}$. Let $\mathcal{C}_{\tilde{S}}$ denote the time instance of the cut point of $\tilde{S}$ (cut time) in this interval and let $\mathcal{C}_{\hat{S}}$ denote the time instance of the cut point of $\hat{S}$ in this interval. Furthermore, since there is no intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$, $\mathcal{C}_{\hat{S}}$ must still be a cut point in the full graph $\hat{S}\cup\tilde{S}$.
Notice that we have $\hat{X}_{n}=\hat{X}[0,\mathcal{C}_{\hat{S}}]+\hat{X}[\mathcal{C}_{\hat{S}},n]\leq\hat{X}[\mathcal{C}_{\hat{S}},n]+n/(\log n)^{2}.$ Similarly, we have $\tilde{X}_{n}\leq\tilde{X}[\mathcal{C}_{\tilde{S}},n]+n/(\log n)^{2}.$ Lastly, we see that $X_{2n}\geq\hat{X}[\mathcal{C}_{\hat{S}},n]+\tilde{X}[\mathcal{C}_{\tilde{S}},n]$. Altogether, we see that on $\Omega^{c}$, we have $E_{n}=\tilde{X}_{n}+\hat{X}_{n}-X_{2n}\leq\frac{2n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$. ∎
@@
Let $E_{n}$ denote the cross term that appears when computing the number of cut points, the effective resistance, or the graph distance. Then, we have,
@@
Let $\Omega$ be the set from Lemma 6.3. Note that by the Cauchy-Schwarz inequality, we have that,
@@ 6.2.1. Our first reduction
The goal of this section is to establish our reduction that showing $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$ is enough to establish $\mathbb{E}[E_{n}^{2}],\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$.
@@
(6.2) $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n}).$
@@
Fix any $\kappa>0$, we will first show that for $n$ large enough, we must have that,
@@
We then have that using the previous expansion on the variance and equation (6.2) that,
$\frac{(2n)^{2}}{(\log(2n))^{f(2n)}}$ $\leq\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\frac{n^{2}}{(\log n)^{f(n)}}$
$+\sqrt{\frac{(\log n)^{-\epsilon}n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}}\sqrt{\frac{2n^{2}}{(\log n)^{f(n)}}}.$
@@
$\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2}}+\frac{n^{2}}{(\log n)^{f(n)+\epsilon}}+C\frac{n^{2}}{(\log n)^{f(n)+\epsilon/2}}+C\frac{n^{2}\sqrt{\log\log n}}{(\log n)^{f(n)/2+1}}$
$\leq\frac{(2+\delta)n^{2}}{(\log n)^{f(n)}}+C\frac{n^{2}\log\log n}{(\log n)^{2}},$
where $\delta>0$ can be set to be any small constant and $C$ is some universal constant (that could depend on $\delta$) and could change from line to line.
@@
In this case, note that for $n$ large enough, we must have that $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{2-\kappa}}$. Thus, we obtain the inequality,
@@
In this case, when $n$ is large enough, we can still obtain the inequality for any small $\delta>0$, $\frac{n^{2}\log\log n}{(\log n)^{2}}\leq\delta\frac{n^{2}}{(\log n)^{f(n)}}$. Thus, we obtain the inequality,
@@
Now, let $n_{0}$ be large enough integer such that both equations (6.3) and (6.4) hold. Furthermore, consider $k$ large enough so that
@@
In this subsection, we will give the estimate only for cut points and the graph distance.
@@
Let $X_{n}$ denote the number of cut points, or the graph distance. Assume that there exists some $\epsilon>0$ such that with probability significantly less than $(\log n)^{-2}$, we have that,
@@
    Notice that we have $E_{n}>l$ implies that $\tilde{X}_{n}\geq\frac{l}{2}$ or $\hat{X}_{n}\geq\frac{l}{2}$. Furthermore, $\Omega$ also holds by Lemma 6.3. Then,
@@
    Furthermore, (again still considering that $E_{n}>l$) in the case $\tilde{X}_{n}\geq\frac{l}{2}$, it must be the case that there is an intersection between $\tilde{S}{[n/(\log n)^{4},n]}$ and $\hat{S}$ by the proof of Lemma 6.3. (Recall, we needed the event $\Omega$ to hold in order to ensure that there is a large reduction in the graph distance.)
@@
We used $\mathbb{E}[\tilde{X}_{n}]\leq\frac{n}{(\log n)^{1/2+o(1)}}$ to show the final inequality. (Note that in [11, Lemma 2.2], they stated that it should be obtained for the graph distance if one follows the proof of [21].) Now, assume that equation (6.5) holds with probability significantly less than $(\log n)^{-2}$. Let $\Omega_{B}$ be the set of walks $\tilde{S}$ such that equation (6.5) holds. We have that,
@@
We start with the following observation. We define a sequence of scales $0=K_{0}<K_{1}<K_{2}<K_{3}<\ldots<K_{f}$ and consider the following bound. Also, let $A(R_{o},R_{i})$ denote the annulus around the origin with outer radius $R_{o}$ and inner radius $R_{i}$. Let $\mathcal{A}_{i,n}:=A[\sqrt{n/K_{i}},\sqrt{n/K_{i+1}}]$.
@@
Eventually, we will choose $K_{i}=\frac{1}{(\log n)^{i(1-\epsilon)}}$ for $i\leq 11$, and after this we choose $K_{i}=2K_{i-1}$ for $i\geq 12$.
Furthermore, we can choose $K_{f}=n^{1/10}$. Note that by the local central limit theorem, the probability a point $\tilde{S}_{k}$ for $k\in[n/(\log n)^{4},n]$ lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{n^{2/10}}{\frac{n^{2}}{(\log n)^{8}}}$. Taking a union bound, the probability that there is a point in $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ that lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{nn^{1/5}(\log n)^{8}}{n^{2}}\ll(\log n)^{-2}$.

### v2

@@ 6.1. Lower bounds for the Cross Term
We will first start by obtaining a lower bound of $\mathbb{E}[E_{n}].$ Our heuristic argument is the following: if there is an intersection between $S[3n/2,2n]$ and $S[0,n]$, then this will form a shortcut in the graph distance between the points $S_{2n}$ and $S_{0}$. The amount that the graph distance would reduce via this shortcut will be related to the graph distance between $S_{n}$ and $S_{2n}$, which, by a law of large numbers, should be $O\left(\frac{n}{(\log n)^{1/2}}\right)$. Furthermore, the probability of the intersection of $S[3n/2,2n]$ and $S[0,n]$ is approximately $O((\log n)^{-1})$. Thus, one would heuristically argue that $\mathbb{E}[E_{n}]=\frac{n}{(\log n)^{3/2\pm o(1)}}$.
However, the exact value of the reduction of the graph distance is a complicated function of the loop intersection structure of the two walks. Thus, knowing that there is a long-range intersection may not be enough to deduce the reduction of the graph distance. Furthermore, to get the exact order $\frac{n}{(\log n)^{3/2}}$, that we desire, we would need to argue that there is an exact independence between the event that there is an intersection between $S[3n/2,2n]$ and $S[0,n]$ and the event that the graph distance between $S_{n}$ and $S_{2n}$ is $O\left(\frac{n}{(\log n)^{1/2}}\right)$. The following lemma will address these issues by finding appropriate events that will allow us to constrain the geometry of the random walk range.
@@
Let $E_{n}$ denote the cross term that appears when considering the number of cut points or the graph distance. We have $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$.
@@
In [21, Section 4], he has already shown $\mathbb{E}[E_{n}]\geq\frac{n}{(\log n)^{3/2+o(1)}}$ for the effective resistance. We obtain the same result with a slightly simpler proof, which derives intersection estimates by using the capacity of the random walk.
@@
$A_{5}=\{\tilde{S}_{i}\leq C_{1}\sqrt{n},\quad\forall i\in[n-n/(\log n)^{\epsilon},n]\}.$
In the above equation, we use the notation $\tilde{X}[a,b]$ to denote the graph distance between the points $\tilde{S}_{a}$ and $\tilde{S}_{b}$ along the graph $\tilde{S}[a,b]$. By [11, Lemma 2.3], $\mathbb{P}(A_{1})=1-o(1)$. Similarly, the law of large numbers ensures that $\mathbb{P}(A_{2})$, $\mathbb{P}(A_{3})=1-o(1)$ (noting [11, Theorem 1.2]) and $\mathbb{P}(A_{4})=1-o(1)$ (see [6, Theorem 1.1]). By the local central limit theorem, we have that $\mathbb{P}(A_{5})=1-\epsilon(C_{1})$, where $\epsilon(C_{1})\to 0$ as $C_{1}\to\infty$. Let $\mathcal{Q}:=A_{1}\cap A_{2}\cap A_{3}\cap A_{4}\cap A_{5}$. From our previous estimates, we can take a union bound to deduce that $\mathbb{P}(\mathcal{Q})\geq 1-2\epsilon(C_{1})$, for $n$ sufficiently large.
Now, we need to define an event on the intersection of $\hat{S}$ with $\tilde{S}$. Let $\tau$ be defined as the first time of intersection between $\hat{S}$ and $\tilde{S}[n-n/(\log n)^{\epsilon},n]$ after time $n-n/(\log n)^{\epsilon}$. That is, $\tau:=\inf\{i>n-n/(\log n)^{\epsilon}:\hat{S}_{i}\in\tilde{S}[n-n/(\log n)^{\epsilon},n]\}$. We define,
@@
On the union $A_{5}\cap B_{2}$, we can apply the local central limit theorem to deduce that $\min_{\hat{a}\in\mathcal{S}}\sum_{i=1}^{n/(\log n)^{\epsilon}}\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\hat{S}_{i+n-n/(\log n)^{\epsilon}}=\hat{a})\geq cn^{-1}(\log n)^{-\epsilon}$, for some constant $c>0$ on $\mathcal{Q}$. Furthermore, the event $A_{5}$ ensures that ${\rm Cap}(\mathcal{S})\geq\frac{1}{2}\frac{\pi^{2}}{8}\frac{n}{(\log n)^{1+\epsilon}}$ on $\mathcal{Q}$. Multiplying these together, we see that there is some $c>0$ such that $\mathbb{P}^{\hat{S}_{n-n/(\log n)^{\epsilon}}}(\tau<n/(\log n)^{\epsilon})>c/(\log n)^{1+2\epsilon}$ on $\mathcal{Q}$. This means that $\mathbb{P}(B_{3}|B_{2}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$. Returning to equation (6.1), we see that we see that we must have $\mathbb{P}(B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q})\geq c/(\log n)^{1+2\epsilon}$ for some $c>0$.
Now, on the event $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$, we claim we have $E_{n}>n/(\log n)^{1/2+\epsilon/2}$. Due to the presence of the cut points from the event $A_{1}$ and the lower bound on the graph distance $\tilde{X}_{n/2}$ from $A_{2}$, we have that $\tilde{X}_{n}\geq\frac{1}{1+\epsilon}\mathbb{E}[X_{n/2}]-\frac{n}{(\log n)^{7/12}}$. By events $A_{3}$, $A_{1}$, and $B_{1}$, the graph distance between the points $\hat{S}_{n}$ and $\tilde{S}_{n}$, is at most $2n/(\log n)^{1/2+\epsilon/2}+2\frac{n}{(\log n)^{7/12}}$. Computing the difference, we now see that $E_{n}>n/(\log n)^{1/2+\epsilon/2}$ on $B_{1}\cap B_{2}\cap B_{3}\cap\mathcal{Q}$ as desired and this event happens with probability $c/(\log n)^{1+2\epsilon}$. By multiplying these two terms together, we see that $\mathbb{E}[E_{n}]>n/(\log n)^{3/2+\epsilon/2+2\epsilon}$. This proof also works even for the number of cut points. ∎
@@
We claim both that $\Omega$ holds with probability $O\left(\frac{\log\log n}{\log n}\right)$ and that on the event $\Omega^{c}$, we must have that the cross term $E_{n}$ for the number of cut points and the graph distance must be less than $\frac{n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$ by Lemma 6.1.
@@
Now, we will check that on $\Omega^{c}$, our cross term can be no more than $\frac{n}{(\log n)^{2}}$. Notice that on $\Omega^{c}$, there must be some cut point in the interval $\left[\frac{n}{(\log n)^{4}},\frac{n}{(\log n)^{2}}\right]$ for both $\tilde{S}$ and $\hat{S}$. Let $\mathcal{C}_{\tilde{S}}$ denote the time instance of the cut point of $\tilde{S}$ (cut time) in this interval, and let $\mathcal{C}_{\hat{S}}$ denote the cut time of $\hat{S}$ in this interval. Furthermore, since there is no intersection between $\hat{S}[n/(\log n)^{4},n]$ and $\tilde{S}$, $\mathcal{C}_{\hat{S}}$ must still be a cut time in the full graph $\hat{S}\cup\tilde{S}$.
Notice that we have $\hat{X}_{n}=\hat{X}[0,\mathcal{C}_{\hat{S}}]+\hat{X}[\mathcal{C}_{\hat{S}},n]\leq\hat{X}[\mathcal{C}_{\hat{S}},n]+n/(\log n)^{2}.$ Similarly, we have $\tilde{X}_{n}\leq\tilde{X}[\mathcal{C}_{\tilde{S}},n]+n/(\log n)^{2}.$ Lastly, $X_{2n}\geq\hat{X}[\mathcal{C}_{\hat{S}},n]+\tilde{X}[\mathcal{C}_{\tilde{S}},n]$. Altogether, on $\Omega^{c}$, $E_{n}=\tilde{X}_{n}+\hat{X}_{n}-X_{2n}\leq\frac{2n}{(\log n)^{2}}\ll\mathbb{E}[E_{n}]$. ∎
@@
Let $E_{n}$ denote the cross term that appears when computing the number of cut points or the graph distance. Then, we have,
@@
Let $\Omega$ be the set from Lemma 6.3. Note that by the Cauchy-Schwarz inequality,
@@ 6.2.1. Our first reduction
The goal of this section is to establish our reduction that showing $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n})$ is sufficient to establish $\mathbb{E}[E_{n}^{2}],\mathrm{Var}(X_{n})\leq\frac{n^{2}}{(\log n)^{2+o(1)}}$.
@@
(6.2) $\mathbb{E}[E_{n}^{2}]\leq\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+(\log n)^{-\epsilon}\mathrm{Var}(X_{n}).$
@@
Fix any $\kappa>0$; we will first show that for $n$ large enough, we must have that,
@@
By using the previous expansion of the variance and equation (6.2), we obtain
$\frac{(2n)^{2}}{(\log(2n))^{f(2n)}}$ $\leq\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+(\log n)^{-\epsilon}\frac{n^{2}}{(\log n)^{f(n)}}$
$+\sqrt{\frac{(\log n)^{-\epsilon}n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}}\sqrt{\frac{2n^{2}}{(\log n)^{f(n)}}}.$
@@
$\frac{2n^{2}}{(\log n)^{f(n)}}+\frac{Cn^{2}\log\log n}{(\log n)^{2+o(1)}}+\frac{n^{2}}{(\log n)^{f(n)+\epsilon}}+C\frac{n^{2}}{(\log n)^{f(n)+\epsilon/2}}+C\frac{n^{2}\sqrt{\log\log n}}{(\log n)^{f(n)/2+1+o(1)}}$
$\leq\frac{(2+\delta)n^{2}}{(\log n)^{f(n)}}+C\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}},$
where $\delta>0$ can be set to be any small constant, and $C$ is some universal constant (that could depend on $\delta$) and could change from line to line.
@@
In this case, note that for $n$ large enough, we must have that $\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}}\leq\delta\frac{n^{2}}{(\log n)^{2-\kappa}}$. Thus, we obtain the inequality,
@@
In this case, when $n$ is large enough, we can still obtain the following inequality for any small $\delta>0$: $\frac{n^{2}\log\log n}{(\log n)^{2+o(1)}}\leq\delta\frac{n^{2}}{(\log n)^{f(n)}}$. Thus, we obtain the inequality,
@@
Now, let $n_{0}$ be a large enough integer such that both equations (6.3) and (6.4) hold. Furthermore, consider $k$ large enough so that
@@
Let $X_{n}$ denote the number of cut points, or the graph distance. Assume that there exists some $\epsilon>0$ such that, with probability significantly less than $(\log n)^{-2}$, we have that,
@@
    Notice that $E_{n}>l$ implies that $\tilde{X}_{n}\geq\frac{l}{2}$ or $\hat{X}_{n}\geq\frac{l}{2}$. Furthermore, $\Omega$ also holds by Lemma 6.3. Then,
@@
    Furthermore, (again still considering that $E_{n}>l$), in the case $\tilde{X}_{n}\geq\frac{l}{2}$, there must be an intersection between $\tilde{S}{[n/(\log n)^{4},n]}$ and $\hat{S}$ by the proof of Lemma 6.3. (Recall, we needed the event $\Omega$ to hold in order to ensure that there is a large reduction in the graph distance.)
@@
[…]
We used $\mathbb{E}[\tilde{X}_{n}]\leq\frac{n}{(\log n)^{1/2+o(1)}}$ to show the final inequality. (Note that in [11, Lemma 2.2], they stated that this bound should be obtained for the graph distance if one follows the proof of [21].) Now, assume that equation (6.5) holds with probability significantly less than $(\log n)^{-2}$. Let $\Omega_{B}$ be the set of walks $\tilde{S}$ such that equation (6.5) holds. We have that,
@@
We start by defining a sequence of scales $0=K_{0}<K_{1}<K_{2}<K_{3}<\ldots<K_{f}$; also, let $A(R_{o},R_{i})$ denote the annulus around the origin with outer radius $R_{o}$ and inner radius $R_{i}$. Let $\mathcal{A}_{i,n}:=A[\sqrt{n/K_{i}},\sqrt{n/K_{i+1}}]$.
@@
Eventually, we will choose $K_{i}=\frac{1}{(\log n)^{i(1-\epsilon)}}$ for $i\leq 11$, and after this, we will choose $K_{i}=2K_{i-1}$ for $i\geq 12$.
Furthermore, we can choose $K_{f}=n^{1/10}$. Note that by the local central limit theorem, the probability that a point $\tilde{S}_{k}$ for $k\in[n/(\log n)^{4},n]$ lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{n^{2/10}}{\frac{n^{2}}{(\log n)^{8}}}$. Taking a union bound, the probability that there is a point in $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ that lies in the ball of radius $n^{1/10}$ around the origin is at most $\frac{nn^{1/5}(\log n)^{8}}{n^{2}}\ll(\log n)^{-2}$.